OpenStudy (anonymous):

Let e = the number of points that Emily scored, m = the number of points that Miguel scored, and v = the number of points that Valerie scored. The three finalists in the talent show are Emily, Miguel, and Valerie. Their combined score is 24. Emily and Miguel's combined score is twice that of Valerie, and Valerie scored only one more point than Miguel. How many points did Emily score? Which equation is needed to represent the situation?

OpenStudy (lochana):

you can make formulas for this V + M + E = 24 2V = M + E V = M + 1

OpenStudy (lochana):

so lets solve it for V ,M and E V + 2V = 24 V = 8

OpenStudy (lochana):

V = M +1 M = 7

OpenStudy (lochana):

V + M + E = 24 E = 24 - M - V E = 24 - 8 - 7 E = 9

OpenStudy (lochana):

so Emily's points is 9